# Model predictive control is almost optimal for restless bandits

### v1

@@ 4.1 First result: LP-update is asymptotic optimal
Our result will show that the LP-update policy is asymptotically optimal as the number of arms $N$ goes to infinity, under an easily verifiable mixing assumption on the transition matrices. To express this condition, for a fixed integer and a sequence of action $\textbf{a}=(a_{1}\dots a_{k})\in\{0,1\}^{k}$, we denote by $P^{\textbf{a}}_{i,j}$ the $(i,j)$th entry of the matrix $\prod_{t=1}^{k}\mathbf{P}^{a_{k}}$. We then denote by $\rho_{k}$ the following quantity [^1] The definition (10) is related to the notion of ergodic coefficient defined in (Puterman, 2014) that used a related constant to prove convergence of value iteration algorithms in span norm for the unconstrained discounted Markov Decision Process.:
@@
In the above equation, the minimum is taken over all possible initial state $s,s^{\prime}$ and all possible sequence of action. The quantity $\rho_{k}$ can be viewed as the probability (under the best coupling) that two arms starting in states $s$ and $s^{\prime}$ reach the same state after $k$ iterations, if the sequence $a_{1}\dots a_{k}$ is used for the first arm while the second arm only uses the action 00. The assumption that we use for our result is that $\rho_{k}>0$ for some integer $k$.
Assumption 1.
@@
While the assumption may look abstract, note that when the $\mathbf{P}^{0}$ matrix is ergodic, it ensures that assumption 1 holds. Indeed in this case, there exists a $k>0$ such that $P^{0\dots 0}_{ij}>0$ for all $i,j$ which would imply that $\rho_{k}>0$. Related assumptions and their relationship to ergodicity can be found in Hernandez-Lerma & Lasserre (2012). Assumption 1 is similar to the unichain and aperiodic condition imposed in Hong et al. (2024b). Note that as this quantity involves the best coupling and not a specific coupling, it is more general than the synchronization assumption used in Hong et al. (2023).
We are now ready to state our first theorem, where we provide a performance bound of the average reward of the LP-update policy, that we denote by $V_{\text{LP},\tau}^{(N)}(\infty)$.
@@
Assume 1. There exist constants $C_{\lambda},C_{\Phi}>0$ such that for any $\epsilon>0$, there exists $\tau(\epsilon)$ such that, Algorithm 1 has the following guarantee of performance:
$V_{\text{LP},\tau(\epsilon)}^{(N)}(\infty)\geq V_{\text{opt}}^{N}(\infty)-\epsilon-\left(\frac{kC_{\Phi}}{\rho}+C_{\lambda}C_{\Phi}+1\right)\frac{(\alpha N-\lfloor{\alpha N}\rfloor)}{N}-\left(\frac{k}{\rho_{k}}+C_{\lambda}\right)\left(\sqrt{\frac{|\mathcal{S}|}{N}}\right)$ (11)
This result shows that the LP-update policy becomes optimal as $\tau$ and $N$ go to infinity. The sub-optimality gap of LP-update decomposes in three terms. The first term corresponds to an upper-bound on the sub-optimality of using the finite-horizon $\tau$ when solving the LP-problem (8). Note, our proof shows that one can take $\tau(\epsilon)=\mathcal{O}(\frac{1}{\epsilon})$. Moreover, in the numerical section, we will show that choosing a small value like $\tau=10$ is sufficient for most problems. The second term corresponds to a rounding error for the budget $\alpha$: for all $\alpha$ such that $N\alpha$ is an integer, this term equals 00. The dominating error term is the last term, $O(1/\sqrt{N})$. It corresponds to the difference between the $N$-arm problem and the LP-problem with $\tau+\infty$.
@@ 4.2 Second result: Exponentially small gap under a stability condition
Theorem 4.1 shows that the sub-optimality gap of the LP-update policy is of order $O(1/\sqrt{N})$ under quite general conditions. While one could not hope for a better convergence rate in general, there are known cases for which one can construct policies that become optimal exponentially fast when $N$ goes to infinity. This is the case for Whittle index under the conditions of indexability, uniform global attractor property (UGAP), non-degeneracy and global exponential stability Gast et al. (2023a). More details can be found in Appendix A. In this section, we show that LP-update also becomes optimal exponentially fast under essentially same conditions as the ones presented in Hong et al. (2024a). The first condition that we impose is that the solution of the above LP-problem is non-degenerate (as defined in Gast et al. (2023a); Hong et al. (2024a)).
Assumption 2 (Non-degenerate).
@@
We assume that the matrix $P^{*}$ is stable, i.e., that all of the norm of all the eigenvalues of $P^{*}$ are strictly smaller than $1$, except for one eigenvalue that equals $1$.
The two above conditions are equivalent to the assumption of non-degeneracy and local stability defined in Hong et al. (2024a).
The last condition that we impose is a technical assumption that simplify the proofs.
Assumption 4 (Unicity).
@@
It guarantees that the LP-update policy is uniquely defined. Note that the assumptions of unicity of the fixed point are often made implicitly in papers when authors talk about “the” optimal solution instead of “an” optimal solution.
@@
Assume 1, 2, 3, and 4 then there exist constants $C^{\prime},C^{\prime\prime}>0$ (independent on $N$ and $\tau$) such that for all $\epsilon>0$, with $\tau(\epsilon)$ (set according to Theorem  4.1) and $N$ such that $\alpha N$ is an integer, Algorithm 1 has the following guarantee of performance
$V_{\text{LP},\tau(\epsilon)}^{(N)}(\infty)\geq V_{\text{opt}}^{N}(\infty)-\epsilon-C^{\prime}e^{-C^{\prime\prime}N}.$ (12)
The first term of the bound above is identical to the one of Theorem 4.1. What is more important is that the last term decays exponentially with $N$.
@@ 5 Proofs: Main Ideas
In this section, we provide the major ingredients of the proofs of the two main theorems. We provide more details for the proof of Theorem 4.1 because this is the more original of the two. The proofs of all lemmas and some details of computation are deferred to Appendix  C.
@@
Three major components are required in-order to complete the proof.
Part 1, Properties of the dynamical control problem (8)
For ${\bm{x}},\bm{u}$, we denote by $\Phi({\bm{x}},\bm{u}):={\bm{x}}\mathbf{P}^{0}+\bm{u}(\mathbf{P}^{1}-\mathbf{P}^{0})$ the deterministic transition kernel, and we recall that the instantaneous reward is $R({\bm{x}},\bm{u}):=\mathbf{r}^{0}\cdot{\bm{x}}+(\mathbf{r}^{1}-\mathbf{r}^{0})$. In Lemma C.2, we establish several properties that relate the finite-horizon problem (8) and the finite-horizon problem (5) that hold under Assumption 1. First we show that the average gain of the finite-time horizon problem (8) converges to the average gain of the infinite-horizon problem, that is $\lim_{\tau\to\infty}W_{\tau}({\bm{x}})/\tau=g^{*}$ for all ${\bm{x}}$. Second, we also show that the bias function $h^{\star}(\cdot):\Delta_{\mathcal{S}}\to\mathbb{R}$ given by :
@@
is well defined and Lipschitz-continuous with constant $k/\rho_{k}$ and that the convergence in (13) is uniform in ${\bm{x}}$. Moreover, the gain and the bias function satisfy
@@
While both these definitions are well known in the average reward unichain MDP setting without constraints, Lemma C.2 establishes these definitions for the constrained, deterministic problem.
@@
Let $({\bm{x}}^{*},\bm{u}^{*})$ be the [^2] For clarity we will present the proof as if this point is unique although the proof can be rendered without this requirement. optimal solution of the infinite-horizon problem (5), and let $l(\textbf{x},\bm{u}):=g^{\star}-\mathbf{r}^{0}\cdot\textbf{x}-(\mathbf{r}^{1}-\mathbf{r}^{0})\cdot\bm{u}$. Following Damm et al. (2014), an optimal control problem with stage cost $l({\bm{x}},\bm{u})$ and dynamic ${\bm{x}}(t+1):=\Phi({\bm{x}},\bm{u})$ is called dissipative if there exists a storage function $\lambda:\Delta_{\mathcal{S}}\to\mathbb{R}$ that satisfy the following equation:
@@
Subject to: ${\bm{x}}(t)$ and $\bm{u}(t)$ satisfy (7) for all $t\in\{0,T-1\}$. (15)
By dissipativity, $L_{\tau}({\bm{x}})$ is monotone increasing. Moreover, we have that $L_{T}(\textbf{x})\geq 0=L_{T}({\bm{x}}^{*})$. Hence, the problem is operated optimally at the fixed point, $(x^{*},u^{*})$. The optimal operation at a fixed point is a key observation, made by several works, (Goldsztajn & Avrachenkov, 2024; Gast et al., 2023b; Yan, 2024; Hong et al., 2023), we recover this result as a natural consequence of dissipativity.
By Lemma C.4, $L_{\tau}(\textbf{x})=\tau g^{\star}-W_{\tau}{({\bm{x}})}+\lambda\cdot{\bm{x}}$ because [^3] In fact, the two control problems are equivalent: $\bm{u}^{*}$ is optimal for (8) if and only if it is for (15). of a telescopic sum of the terms $\lambda\cdot{\bm{x}}(t)$. Combining this with (13) implies that $\lim_{\tau\to\infty}L_{\tau}({\bm{x}})=h^{\star}({\bm{x}})+\lambda\cdot{\bm{x}}$. As $L_{\tau}({\bm{x}})$ is monotone, it follows that, for any $\epsilon>0$, there exists a $\tau(\epsilon)$ such that for $L_{\tau(\epsilon)}({\bm{x}})-L_{\tau(\epsilon)-1}({\bm{x}})<\epsilon$. Note that the fact that $L_{\tau}({\bm{x}})$ is monotone is crucial to obtain this property, leading us to use dissipativity. Putting the three components together, we can now prove the final steps.
Proof of Theorem 4.1.
We begin by considering the difference between the optimal value and the model predictive value function, we let ${\bm{X}(t)}$ denote the empirical distribution of the arms at time $t$ and $\bm{U}(t)$ be the corresponding empirical distribution of the actions. We drop the superscript $N$ for convenience in the proof below, but it should be noted that $\bm{U}(t)$ is always obtained from a randomized rounding procedure and hence, dependent on $N$.
$V_{\text{opt}}^{N}(\textbf{x})-V_{\text{LP},T}^{(N)}(\textbf{x})$ $\leq\lim_{T\to\infty}\frac{1}{\tau}\sum_{t=0}^{T-1}\mathbb{E}\left[g^{\star}-R({\bm{X}}(t),{\bm{U}}(t))\right]$
@@
The first inequality follows from the well known result $V_{\text{opt}}^{N}(\textbf{x})\leq g^{\star}$, (Yan, 2024; Gast et al., 2023a; Hong et al., 2023). The last inequality follows from randomized rounding and Lemma B.1. Let $(A):=\lim_{T\to\infty}\frac{1}{\tau}\sum_{t=0}^{T-1}\mathbb{E}\left[l({\bm{X}}(t),\bm{u}(t))\right]$ denotes the first term of (16). Adding and subtracting the storage cost we have:
@@
$L_{\tau}({\bm{x}})-L_{\tau-1}(\Phi({\bm{x}},\bm{u}))=L_{\tau}({\bm{x}})-L_{\tau}(\Phi({\bm{x}},\bm{u}))+L_{\tau}(\Phi({\bm{x}},\bm{u}))-L_{\tau-1}(\Phi({\bm{x}},\bm{u})).$
By choosing $\tau=\tau(\epsilon)$ so that, $L_{\tau}({\bm{x}})-L_{\tau-1}({\bm{x}})<\epsilon$, we can ensure,
$L_{\tau}({\bm{x}})-L_{\tau-1}(\Phi({\bm{x}},\bm{u}))\leq L_{\tau}({\bm{x}})-L_{\tau}(\Phi({\bm{x}},\bm{u}))+\epsilon$ (17)
Plugging these inequalities together, introducing a telescopic sum, and manipulating the order of variables slightly (shown in Appendix C.4), we have the following relation:
$(A)\leq\epsilon+\lim_{T\to\infty}$ $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[L_{\tau}({\bm{X}}(t+1))-L_{\tau}(\Phi({\bm{X}}(t),\bm{u}(t)))-\lambda\cdot{\bm{X}}(t+1)+\lambda\cdot\Phi({\bm{X}}(t),\bm{u}(t))\right].$ (18)
$L_{\tau}(\textbf{x})$ is Lipschitz in x for any $\tau$. Since, $\mathbb{E}[{\bm{X}}(t+1)|{\bm{X}}(t),\bm{U}(t)]=\Phi({\bm{X}}(t),\bm{U}(t))$, Lemma 1 of Gast et al. (2023b) implies:
$\|\mathbb{E}[{\bm{X}}(t+1)|{\bm{X}}(t),\bm{U}(t)]-\Phi({\bm{X}}(t),\bm{u}(t))\|_{1}\leq\frac{\sqrt{|\mathcal{S}|}}{\sqrt{N}}+\frac{C_{\Phi}(\alpha N-\lfloor{\alpha N}\rfloor)}{N},$
where $C_{\Phi}$ is the Lipschitz constant of the map $\Phi$. The Lipschitz constant of $L_{\tau}(\cdot)$ in the $l1$ norm is $\frac{k}{\rho_{k}}$ and $C_{\lambda}$ is equal to $\|\lambda\|_{\infty}$. Passing to the limit $\tau\to\infty$ we have:
$V_{\text{opt}}^{N}(\textbf{x})-V_{\text{LP},\tau}^{(N)}(\textbf{x})\leq\epsilon+\left(\frac{k}{\rho}+C_{\lambda}\right)\left(\frac{\sqrt{|\mathcal{S}|}}{\sqrt{N}}+\frac{C_{\Phi}(\alpha N-\lfloor{\alpha N}\rfloor)}{N}\right)+\frac{(\alpha N-\lfloor{\alpha N}\rfloor)}{N}$ (19)

### v2

@@ 4.1 First result: LP-update is asymptotic optimal
Our result will show that the LP-update policy is asymptotically optimal as the number of arms $N$ goes to infinity, under an easily verifiable mixing assumption on the transition matrices. To express this condition, for a fixed integer and a sequence of actions $\textbf{a}=(a_{1}\dots a_{k})\in\{0,1\}^{k}$, we denote by $P^{\textbf{a}}_{i,j}$ the $(i,j)$th entry of the matrix $\prod_{t=1}^{k}\mathbf{P}^{a_{k}}$. We then denote by $\rho_{k}$ the following quantity [^1] The definition (10) is related to the notion of ergodic coefficient defined in (Puterman, 2014) that used a related constant to prove convergence of value iteration algorithms in span norm for the unconstrained discounted Markov Decision Process.:
@@
In the above equation, the minimum is taken over all possible initial states $s,s^{\prime}$ and all possible sequence of actions. The quantity $\rho_{k}$ can be viewed as the probability (under the best coupling) that two arms starting in states $s$ and $s^{\prime}$ reach the same state after $k$ iterations, if the sequence $a_{1}\dots a_{k}$ is used for the first arm while the second arm only uses the action 00. The assumption that we use for our result is that $\rho_{k}>0$ for some integer $k$.
Assumption 1
@@
While the assumption may look abstract, note that when the $\mathbf{P}^{0}$ matrix is ergodic, it ensures that assumption 1 holds. Indeed in this case, there exists a $k>0$ such that $P^{0\dots 0}_{ij}>0$ for all $i,j$ which would imply that $\rho_{k}>0$. Related assumptions and their relationship to Ergodicity can be found in Hernandez-Lerma and Lasserre (2012). Assumption 1 is similar to the unichain and aperiodic condition imposed in Hong et al. (2024b). Note that this quantity involves the best coupling and not a specific coupling; it is more general than the synchronization assumption from Hong et al. (2023).
We are now ready to state our first theorem, in which we provide a performance bound of the average reward of the LP-update policy, that we denote by $V_{\text{LP -}\tau}^{(N)}$.
@@
Assume 1 with ergodicity constant $\rho_{k}$ for some fixed integer $k$. Fix a positive constant $\epsilon>0$; then there exists $\tau(\epsilon)$ such that, Algorithm 1 has the following guarantee of performance:
$V_{\text{LP -}\tau(\epsilon)}^{(N)}\geq V_{\text{opt}}^{\infty}-2\epsilon-\left(\frac{2k}{\rho_{k}}\left[3+\frac{2k\alpha}{\rho_{k}}\right]+1\right)\frac{\alpha N-\lfloor{\alpha N}\rfloor}{N}-\frac{k}{\rho_{k}}\left(3+\frac{2k\alpha}{\rho_{k}}\right)\sqrt{\frac{|\mathcal{S}|}{N}}$ (11)
As $V_{\text{opt}}^{N}\leq V_{\text{opt}}^{\infty}$, this result shows that the LP-update policy becomes optimal as $\tau$ and $N$ go to infinity. The sub-optimality gap of LP-update decomposes in three terms. The first term corresponds to an upper-bound on the sub-optimality of using the finite-horizon $\tau$ when solving the LP-problem (8). Our proof shows that one can take $\tau(\epsilon)=\mathcal{O}(\frac{1}{\epsilon})$, and in the numerical section, we will show that choosing a small value like $\tau=10$ is sufficient for most problems. The second term corresponds to a rounding error: for all $\alpha$ such that $N\alpha$ is an integer, this term equals 00. The dominating error term is the last term, $O(1/\sqrt{N})$. It corresponds to the difference between the $N$-arm problem and the LP-problem with $\tau=+\infty$.
@@ 4.2 Second result: Exponentially small gap under a stability condition
Theorem 4.1 shows that the sub-optimality gap of the LP-update policy is of order $O(1/\sqrt{N})$ under quite general conditions. While one could not hope for a better convergence rate in general, there are known cases for which one can construct policies that become optimal exponentially fast when $N$ goes to infinity. This is the case for Whittle index under the conditions of indexability, uniform global attractor property (UGAP), non-degeneracy and global exponential stability Gast et al. (2023a). More details can be found in Appendix A. In this section, we show that LP-update also becomes optimal exponentially fast under essentially the same conditions as the ones presented in Hong et al. (2024a). The first condition that we impose is that the solution of the above LP-problem is non-degenerate (as defined in Gast et al. (2023a); Hong et al. (2024a)).
Assumption 2 (Non-degenerate)
@@
We assume that the matrix $\mathbf{P}^{*}$ is stable, i.e., that the $l_{2}$ norm of all but one of the Eigenvalues of $\mathbf{P}^{*}$ are strictly smaller than 1.
Both these conditions are equivalent to the assumption of non-degeneracy and local stability defined in Hong et al. (2024a).
The last condition that we impose is a technical assumption that simplifies the proofs.
Assumption 4 (Unicity)
@@
This assumption guarantees that the LP-update policy is uniquely defined. Note that the assumptions of unicity of the fixed point are often made implicitly in papers when authors talk about “the” optimal solution instead of “an” optimal solution. The assumptions of non-degeneracy and unicity are not restrictive since, degenerate solutions essentially occupy a zero-measure set: if a problem is degenerate (or has multiple solutions), then adding a small noise to the parameters will make this problem non-degenerate (or will guarantee the uniqueness of the solution). The situation is, however, different for the local stability assumption: if Assumption 3 does not hold for a given problem, then it will not hold for any problem that is close enough to this problem. Moreover, we believe that this assumption is necessary to derive a tighter bound in terms of the number of arms (Hong et al. (2024a)).
@@
Assume 1, 2, 3, and 4; then there exist constants $C^{\prime},C^{\prime\prime}>0$ (independent of $N$ and $\tau$) such that for all $\epsilon>0$, with $\tau(\epsilon)$ (set according to Theorem  4.1) and $N$ such that $\alpha N$ is an integer, Algorithm 1 has the following guarantee of performance:
$V_{\text{LP -}\tau(\epsilon)}^{(N)}\geq V_{\text{opt}}^{\infty}-2\epsilon-C^{\prime}e^{-C^{\prime\prime}N}.$ (12)
The first term of the bound above is identical to the one used in Theorem 4.1. What is more important is that the last term decays exponentially with $N$. The results of Theorem 4.1 meet, order-wise the best known lower bounds without local stability conditions from Hong et al. (2024a) whereas Theorem 4.2 meets the best exponential bounds with stability conditions. Our theorem shows that the LP-update policy also benefit from similar performance guarantees (while performing better for small values of $N$). Previous results on LP-update policies were only able to match our bounds under the UGAP assumptions. Our relaxation of these assumptions can be regarded as one of the main contributions of our works.
The bounds that we obtain measure the performance gap $V_{\text{LP -}\tau(\epsilon)}^{(N)}-V_{\text{opt}}^{\infty}$ which is the difference between the value of the LP-update policy for the system of size $N$ and the value of the relaxed LP. This performance gap is known to be of order at least $\Omega(1/\sqrt{N})$ for general degenerate problems and at least $\exp{-O{N}}$ for general non-degenerate problems. As $V_{\text{opt}}^{\infty}\geq V_{\text{opt}}^{N}\geq$, this implies that the same bounds of $O(1/\sqrt{N})$ and $\exp(-\Omega(N))$ hold for the sub-optimality gap $V_{\text{LP -}\tau(\epsilon)}^{(N)}-V_{\text{opt}}^{N}$. In a recent paper, Yan et al. (2024) directly study this sub-optimality gap directly without comparing the benchmark to $V_{\text{opt}}^{\infty}$ for the finite horizon problem.
@@ 5 Proofs: Main Ideas
In this section, we provide the major ingredients of the proofs of the two main theorems. We provide more details for the proof of Theorem 4.1 because this is the more original of the two. The proofs of all lemmas and some details of computation are deferred to the supplementary material.
@@
Three major components are required in order to complete the proof.
Part 1, Properties of the dynamical control problem (8):
For ${\bm{x}},\bm{u}$, we denote by $\Phi({\bm{x}},\bm{u}):={\bm{x}}\mathbf{P}^{0}+\bm{u}(\mathbf{P}^{1}-\mathbf{P}^{0})$ the deterministic transition kernel, and we recall that the instantaneous reward is $R({\bm{x}},\bm{u}):=\mathbf{r}^{0}\cdot{\bm{x}}+(\mathbf{r}^{1}-\mathbf{r}^{0})$. In Lemma C.2, we establish several properties that relate the finite-horizon problem (8) and the finite-horizon problem (5) that hold under Assumption 1. First, we show that the average gain of the finite-time horizon problem (8) converges to the average gain of the infinite-horizon problem, that is $\lim_{\tau\to\infty}W_{\tau}({\bm{x}})/\tau=g^{*}$ for all ${\bm{x}}$. Second, we also show that the bias function $h^{\star}(\cdot):\Delta_{\mathcal{S}}\to\mathbb{R}$ given by :
@@
is well defined and Lipschitz-continuous with constant $k/\rho_{k}$ and the convergence in (13) is uniform in ${\bm{x}}$. Moreover, the gain and the bias function satisfy
@@
While both these definitions are well known in the average reward unichain MDP, single arm setting without constraints, we establish these definitions for the constrained, deterministic problem. Note the difference in the second term on the right hand side of the fixed point equation: typically the right hand side in the single arm problem takes the form $R(s,a)+\sum_{s^{\prime}}P(s^{\prime}|s,a)h(s^{\prime})$, whereas here the expectation is inside rather than outside!
@@
Let $({\bm{x}}^{*},\bm{u}^{*})$ be the [^2] For clarity we will present the proof as if this point is unique although the proof also holds without this requirement. optimal solution of the infinite-horizon problem (5), and let $l(\textbf{x},\bm{u}):=g^{\star}-\mathbf{r}^{0}\cdot\textbf{x}-(\mathbf{r}^{1}-\mathbf{r}^{0})\cdot\bm{u}$. Following Damm et al. (2014), an optimal control problem with stage cost $l({\bm{x}},\bm{u})$ and dynamic ${\bm{x}}(t+1):=\Phi({\bm{x}},\bm{u})$ is called dissipative if there exists a storage function $\lambda:\Delta_{\mathcal{S}}\to\mathbb{R}$ that satisfies the following equation:
@@
Subject to: ${\bm{x}}(t)$ and $\bm{u}(t)$ satisfy (7) for all $t\in\{0,\tau-1\}$. (15)
By dissipativity, $L_{\tau}({\bm{x}})$ is monotone increasing. Moreover, we have that $L_{\tau}(\textbf{x})\geq 0=L_{\tau}({\bm{x}}^{*})$. Hence, the problem is operated optimally at the fixed point, $(x^{*},u^{*})$. The optimal operation at a fixed point is a key observation, made by several works, (Goldsztajn and Avrachenkov, 2024; Gast et al., 2023b; Yan, 2024; Hong et al., 2023), we recover this result as a natural consequence of dissipativity. By Lemma C.6, $L_{\tau}(\textbf{x})=\tau g^{\star}-W_{\tau}{({\bm{x}})}+\lambda\cdot{\bm{x}}$ because [^3] In fact, the two control problems are equivalent: $\bm{u}^{*}$ is optimal for (8) if and only if it is for (15). of a telescopic sum of the terms $\lambda\cdot{\bm{x}}(t)$. Combining this with (13) implies that $\lim_{\tau\to\infty}L_{\tau}({\bm{x}})=h^{\star}({\bm{x}})+\lambda\cdot{\bm{x}}$. We will denote this limit by $L_{\infty}({\bm{x}})$. As $L_{\tau}({\bm{x}})$ is monotone, it follows that, for any $\epsilon>0$, there exists a $\tau(\epsilon)$ such that $|L_{\infty}({\bm{x}})-L_{\tau(\epsilon)}({\bm{x}})|<\epsilon$. Note, the monotonicity of $L_{\tau}({\bm{x}})$ is crucial to obtain this property, leading us to use dissipativity. Putting the components together, we can now prove the final steps.
Proof of Theorem 4.1
We begin by considering the difference between the optimal value and the model predictive value function. We let ${\bm{X}(t)}$ denote the empirical distribution of the arms at time $t$ and $\bm{U}(t)$ be the corresponding empirical distribution of the actions. We drop the superscript $N$ for convenience in the proof below, but it should be noted that $\bm{U}(t)$ is always obtained from a randomized rounding procedure and hence, is dependent on $N$.
$V_{\text{opt}}^{N}-V_{\text{LP -}\tau}^{(N)}$ $\leq\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[g^{\star}-R({\bm{X}}(t),{\bm{U}}(t))\right]$
@@
The first inequality follows from the well known result $V_{\text{opt}}^{N}\leq g^{\star}$, (Yan, 2024; Gast et al., 2023a; Hong et al., 2023). The last inequality follows from randomized rounding and Lemma B.1. Let $(A):=\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[l({\bm{X}}(t),\bm{u}(t))\right]$ denote the first term of (16). Adding and subtracting the storage cost we have:
@@
$L_{\tau}({\bm{x}})-L_{\tau-1}(\Phi({\bm{x}},\bm{u}))=L_{\tau}({\bm{x}})-L_{\infty}(\Phi({\bm{x}},\bm{u}))+L_{\infty}(\Phi({\bm{x}},\bm{u}))-L_{\tau-1}(\Phi({\bm{x}},\bm{u})).$
By choosing $\tau=\tau(\epsilon)$ we have, $\max\{|L_{\tau-1}({\bm{x}})-L_{\infty}({\bm{x}})|,|L_{\tau}({\bm{x}})-L_{\infty}({\bm{x}})|\}<\epsilon$. Plugging these inequalities together, introducing a telescopic sum, and manipulating the order of variables slightly (shown in Appendix  C.4), the term $(A)$ is smaller than
$2\epsilon+\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[L_{\infty}({\bm{X}}(t+1))-L_{\infty}(\Phi({\bm{X}}(t),\bm{u}(t)))-\lambda\cdot[{\bm{X}}(t{+}1){-}\Phi({\bm{X}}(t),\bm{u}(t))]\right].$ (17)
By Lemma 1 of Gast et al. (2023b), we have $\|\mathbb{E}[{\bm{X}}(t{+}1)|{\bm{X}}(t),\bm{U}(t)]-\Phi({\bm{X}}(t),\bm{U}(t))\|_{1}\leq\sqrt{|S|}/\sqrt{N}$. Moreover, our rounding procedure implies that $\|\Phi({\bm{X}}(t),\bm{U}(t))-\Phi({\bm{X}}(t),\bm{u}(t))\|_{1}\leq C_{\Phi}(\alpha N-\lfloor\alpha N\rfloor)/N$ where $C_{\Phi}\leq 2$ is the Lipschitz-constant of the map $\Phi$ (see C.5). With a few extra steps (see C.5)we get:
$(A)\leq 2\epsilon+(C_{L}+\|\lambda\|_{\infty})\left(\frac{\sqrt{|S|}}{\sqrt{N}}+2\frac{\alpha N-\lfloor\alpha N\rfloor}{N}\right),$ (18)
where $C_{L}$ is a Lipschitz constant for the limiting map $L_{\infty}(\cdot)$ and $\|\lambda\|_{\infty}$ is the infinity-norm of the Lagrange multiplier $\lambda$. In Appendix Appendix C.3[Lemma C.6], we show that $C_{L}$ is bounded above by $k/\rho_{k}+\|\lambda\|_{\infty}$, further, in Appendix F we show that $\|\lambda\|_{\infty}\leq(k/\rho_{k})(1+\alpha k/\rho_{k})$. The theorem follows by substituting these values into (18) and adding a term $(\alpha N-\lfloor{\alpha N}\rfloor)/N$ coming from (16). See Appendix  C.5.
@@
Refer to caption Refer to caption Refer to caption
(a) Random example (b) Example Hong et al. (2023) (c) Example Yan (2022)
Figure 1: Performance as a function of $N$